OpenStudy (anonymous):

Solve the equation by factoring. (Do not use mixed numbers in your answers. 5t^2+8=13t

OpenStudy (anonymous):

\[5t^2 - 13t + 8 = 0\] \[(5t - 8)(5t - 5) = 0\] \[(5t - 8)(t - 1) = 0\] \[t = \frac{8}{5} \space\vee\space t = 1\]